OpenStudy (zzr0ck3r):

Abstract algebra given finite ring R with unity 1 not equal to 0, and R has no zero divisors. prove R is a division ring.

OpenStudy (zzr0ck3r):

so if R is finite, and has n elements, then I can take some element \(a\ne0\) in ring and take the left product with and each of the n-1 non zero element, and get back n-1 elements(each will be different because we have cancellation). So since we get back n-1 elements one of them must be 1 So I have shown there exists \(a_1\) st \(a*a_1=1\) now we can do the same thing again to show there is some \(a_2 \ s.t. \ a_2*a=1\) how do i show that \(a_1=a_2\)?

OpenStudy (zzr0ck3r):

@oldrin.bataku when you get a sec, could you look at this. please sir

OpenStudy (zzr0ck3r):

@ChristopherToni that goes for you as well:)

OpenStudy (anonymous):

let me see

OpenStudy (zzr0ck3r):

every place I see this proof, they stop here. no one ever shows that the left and right elemets are the same...

OpenStudy (anonymous):

>we have cancellation what do you mean here?

OpenStudy (zzr0ck3r):

ab=cb imples a=c

OpenStudy (zzr0ck3r):

we get that for free because there are no zero divisors

OpenStudy (zzr0ck3r):

sorry i forgot to add that to the question.

OpenStudy (zzr0ck3r):

I edited it.

OpenStudy (anonymous):

ah there are no zero-divisors... good good

OpenStudy (zzr0ck3r):

sorry about that

OpenStudy (anonymous):

ok got it

OpenStudy (anonymous):

\(a_1=1\cdot a_1=(a_2a)a_1=a_2(aa_1)=a_2\cdot1=a_2\)

OpenStudy (zzr0ck3r):

ahh lol, wow. I thought I needed to some up with a bijection from R to R and do some ugly stuff. ty again:)

OpenStudy (zzr0ck3r):

some = come

OpenStudy (anonymous):

NP :-p glad to help

OpenStudy (zzr0ck3r):

impressive as always

OpenStudy (anonymous):

do you know how long it takes to reach \(90\) from \(89\) SmartScore-wise? it is automatic, yeah?